A chemist needs to make 100 mL of a 30% alcohol solution by mixing 25% and 55% solutions. How much of the 55% solution should the chemist use?

Answer by TimothyLamb(4379)

---
25x + 55y = 30*100
x + y = 100
---
put the system of linear equations into standard form
---
25x + 55y = 3000
x + y = 100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 83.333333 ml of 25%
y= 16.666666 ml of 55%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php